# A Survey on LLM-as-a-Judge

### v2

@@ 6.2.3. AI for Science
LLMs have demonstrated notable potential in scientific fields, especially in areas like medical question-answering and mathematical reasoning, where they serve as evaluators to improve accuracy and consistency. In medical applications, studies by Brake et al. (2024) and Krolik et al. (2024) showed that models like LLaMA2 can assess clinical notes and Q&A responses with a level of accuracy approaching that of human experts (Brake and Schaaf, 2024; Krolik et al., 2024). This approach leverages prompt engineering to embed expert knowledge, enabling LLMs to handle complex, nuanced information, which provides a reliable first-line assessment that lessens the load on human experts.
In mathematical reasoning, reinforcement learning (RL) and cooperative reasoning methods further enhance LLM’s capability as an evaluator, especially for theory-proofing works (Lu et al., 2023). For example, WizardMath was introduced by employing RL through step-by-step feedback to refine reasoning in mathematical tasks (Luo et al., 2023b). Zhu et al. (2023) proposed a Cooperative Reasoning (CoRe) framework that combines generation and verification to mimic human-like dual-process reasoning, enhancing the model’s problem-solving accuracy (Zhu et al., 2023c). Additionally, Lu et al. (2023) developed MathVista, a benchmark for evaluating mathematical reasoning in visual contexts, which assesses LLMs like GPT-4V on tasks involving mathematical reasoning with visual components (Lu et al., 2023). These methods highlight the value of combining RL, cooperative reasoning, and prompt engineering in improving LLMs’ evaluative and reasoning skills across mathematical reasoning.
@@ 6.2.4. Others
LLMs have also been employed as evaluators to enhance efficiency and consistency across various fields. In software engineering, a method was proposed for using LLMs to evaluate bug report summarizations, demonstrating high accuracy in assessing correctness and completeness, even surpassing human evaluators who experienced fatigue (Kumar et al., 2024a). This approach offers a scalable solution for evaluation. In education, automated essay scoring and revising were explored using open-source LLMs, achieving performance comparable to traditional deep-learning models. Techniques such as few-shot learning and prompt tuning improved scoring accuracy, while revisions effectively enhanced essay quality without compromising original meaning (Song et al., 2024b). In content moderation, an LLM-based approach was developed to identify rule violations on platforms like Reddit, achieving high true-negative rates but encountering challenges with complex rule interpretation, emphasizing the necessity of human oversight for nuanced cases (Kolla et al., 2024). In behavioral sciences, the LLM-as-a-Judge framework was evaluated for assessing user preferences based on personas, revealing limitations in reliability and consistency due to oversimplified personas, but improved significantly through verbal uncertainty estimation, achieving high agreement with human evaluations for high-certainty cases (Dong et al., 2024). These applications of LLMs as evaluators highlight their growing potential in diverse sectors, emphasizing the need for integrating domain-specific knowledge and refining methodologies.
Moreover, LLMs as evaluators demonstrate significant advantages in qualitative assessments that are difficult to quantify, such as evaluating service quality, analyzing user experience feedback, and assessing creative content like art or literature reviews. LLMs’ capability to understand and generate nuanced language makes them well-suited for subjective evaluation tasks traditionally requiring human judgment. Future research will increasingly focus on these areas, exploring how LLMs as judges can enhance assessment accuracy and consistency where traditional quantitative methods fall short.
@@ 7.1. Reliability
Another significant aspect of reliability is fairness and generalization. Evaluations by LLM-as-a-judge can exhibit considerable inconsistency depending on the context. This is why prompt-based methods are often used to improve LLM-as-a-judge performance. However, challenges related to fairness and generalization may arise due to the sensitivity of prompt engineering.
For example, the order of the examples in the context can significantly affect the model’s output, leading to unfair evaluations if the examples are poorly arranged. Moreover, LLMs struggle to handle long context windows effectively, often showing degraded performance or prioritizing later examples in the sequence. These issues raise concerns about fairness and generalization in LLM-based evaluations.
@@ 7.2. Robustness
7.3. Powerful Backone Model
@@ 8.1. More Reliable LLM-as-a-Judge
inally, the uncertain and evolving nature of robustness risks underscores the necessity of proactive mitigation strategies. These strategies should include the development of adversarial training techniques tailored to judgment tasks, the integration of robust uncertainty quantification methods, and the implementation of human-in-the-loop systems to oversee critical decisions. By addressing these challenges, we can build more resilient and dependable systems capable of maintaining high levels of reliability even under adversarial conditions.
@@ 8.2. LLM-as-a-Judge for Data Annotation
Despite its wide applications, data annotation poses significant challenges for current machine learning models due to the complexity, subjectivity, and diversity of data. This process requires domain expertise and is resource-intensive, particularly when manually labeling large datasets. Advanced LLMs such as GPT-4 (OpenAI, 2023a), Gemini (Google, 2023), and LLaMA-2 (Touvron et al., 2023c) offer a promising opportunity to revolutionize data annotation. LLMs serve as more than just tools but play a crucial role in improving the effectiveness and precision of data annotation. Their ability to automate annotation tasks (Zhang et al., 2023e), ensure consistency across large volumes of data, and adapt through fine-tuning or prompting for specific domains (Song et al., 2024a), significantly mitigates the challenges encountered with traditional annotation methods, setting a new standard for what is achievable in the realm of NLP.
Whether in the field of scientific research or industry, we are all still suffering from insufficient target data and domain-specific data, or situations where the data quality is not high enough. Assuming that LLM-as-a-judge can achieve stable performance and be fair and reliable, we can use LLM to annotate data in scenarios where data is insufficient to expand the data. In scenarios with low data quality, we can assess the data quality through LLM, and label the quality tags to achieve the goal of selecting high-quality data. Currently, we have not been able to experimentally rely solely on LLM for a reliable evaluation of various different scenarios of data; most of the time, we still rely on human annotation to ensure professionalism and reliability. LLM-as-a-judge often needs to learn from human annotations in order to perform certain labeling tasks.
@@ 8.3. MLLM-as-a-Judge
AI systems are evolving into highly versatile and multifunctional entities. Traditionally, specialized models were required for distinct language processing tasks, such as sentiment analysis, syntactic parsing, and dialogue modeling. However, large language models (LLMs) have demonstrated competence across these tasks using a single set of weights (Srivastava et al., 2022). Similarly, advancements are being made toward unified systems capable of processing multiple data modalities. Instead of employing distinct architectures for processing text, audio, and images, recent models like GPT-4o (OpenAI, 2023b), Gemini (Google, 2023), and LLaVA (Liu et al., 2023c) integrate these capabilities within a single framework. These developments highlight a growing trend toward unification in the structure and functionality of AI systems, which extends to the emerging paradigm of LLM-as-a-Judge.
Currently, the emergence of MLLM-as-a-Judge frameworks (Chen et al., 2024c) for evaluating models can be observed. However, research exploring how MLLM-as-a-Judge could be applied to the evaluation of data or agents remains limited. Beyond model evaluation, MLLM-as-a-Judge, much like LLM-as-a-Judge, is envisioned to have the capability to assess or annotate data, function as a Reward Model, or serve as a Verifier within intermediate reasoning processes. These expanded roles would allow MLLM-as-a-Judge to contribute more broadly to the AI pipeline.
The future of evaluation lies in developing robust multimodal evaluators capable of reasoning about and assessing complex content spanning text, audio, images, and video. While current multimodal LLMs exhibit promising capabilities, they often lack the reasoning depth and reliability of their text-based counterparts. Future research must address these limitations, with a focus on enhancing reasoning capabilities, improving reliability, and enabling seamless integration across modalities. A practical multimodal evaluator has the potential to not only advance AI research but also enable new applications in areas such as multimodal content moderation and automated knowledge extraction.
@@ 8.5. LLM-as-a-Judge for LLM Optimization
Despite these encouraging developments, current research efforts are still in their infancy. Future work should focus on broadening the application domains and strategies for implementing LLM-as-a-Judge, especially in complex, multimodal scenarios. Furthermore, a systematic assessment of its reliability and generalization capabilities will be critical for fully realizing the potential of LLM-as-a-Judge in enhancing model performance and robustness.
@@ 9. Conclusion
LLM-as-a-Judge represents an LLM-based paradigm for evaluation, offering a compelling alternative to traditional methods that rely either on human experts or quantitative assessments. By leveraging the powerful capabilities of LLMs, this framework excels in scoring, rating, and evaluating tasks with exceptional scalability, adaptability, and precision. These qualities align directly with the growing demand for efficient and robust evaluation systems across diverse fields.
However, fully realizing the potential of LLM-as-a-Judge requires addressing significant challenges, particularly those concerning reliability. The evaluation process must account for factors such as consistency, bias mitigation, and contextual adaptability, ensuring that the system operates under well-defined rules while delivering human-aligned and objective results. Each stage of development—from dataset curation and model fine-tuning to evaluation protocol standardization—demands careful consideration to uphold reliability. Simply deploying LLM-as-a-Judge does not guarantee outcomes that are both accurate and aligned with evaluation needs. Building reliable systems requires systematic efforts to evaluate and refine its outputs, while addressing gaps in definitions, practices, and research.
This survey has provided a comprehensive review of the LLM-as-a-Judge landscape, offering insights into its current state and practical strategies for improvement. By addressing these challenges through robust methodologies and innovative benchmarks, LLM-as-a-Judge can evolve into a highly reliable and practical tool for diverse applications. Ultimately, while the current capabilities of LLMs are impressive, their effectiveness as evaluators depends on thoughtful design and systematic refinement. We hope this work serves as a foundation for further research, guiding the development of more reliable LLM-as-a-Judge systems and facilitating their widespread deployment across academic, industrial, and societal domains.

### v3

@@ 6.2.3. AI for Science
LLMs have demonstrated notable potential in scientific fields, especially in areas like medical question-answering and mathematical reasoning, where they serve as evaluators to improve accuracy and consistency. In the medical field, studies by Brake et al. (2024) and Krolik et al. (2024) showed that models like LLaMA2 can assess clinical notes and Q&A responses with a level of accuracy approaching that of human experts (Brake and Schaaf, 2024; Krolik et al., 2024). This approach leverages prompt engineering to embed expert knowledge, enabling LLMs to handle complex, nuanced information, which provides a reliable first-line assessment that lessens the load on human experts.
In mathematical reasoning, reinforcement learning (RL) and cooperative reasoning methods further enhance LLM’s capability as an evaluator, especially for theorem-proving tasks (Lu et al., 2023). For example, WizardMath was introduced by employing RL through step-by-step feedback to refine reasoning in mathematical tasks (Luo et al., 2023b). Zhu et al. (2023) proposed a Cooperative Reasoning (CoRe) framework that combines generation and verification to mimic human-like dual-process reasoning, enhancing the model’s problem-solving accuracy (Zhu et al., 2023c). Additionally, Lu et al. (2023) developed MathVista, a benchmark for evaluating mathematical reasoning in visual contexts, which assesses LLMs like GPT-4V on tasks involving mathematical reasoning with visual components (Lu et al., 2023). These methods highlight the value of combining RL, cooperative reasoning, and prompt engineering in improving LLMs’ evaluative and reasoning skills across mathematical reasoning.
@@ 6.2.4. Others
LLMs have also been employed as evaluators to enhance efficiency and consistency across various fields. In software engineering, a method was proposed for using LLMs to evaluate bug report summarizations, demonstrating high accuracy in assessing correctness and completeness, even surpassing human evaluators who experienced fatigue (Kumar et al., 2024a). This approach offers a scalable solution for evaluation. In education, automated essay scoring and revising have been explored using open-source LLMs, achieving performance comparable to traditional deep-learning models. Techniques such as few-shot learning and prompt tuning improved scoring accuracy, while revisions effectively enhanced essay quality without compromising original meaning (Song et al., 2024b). In content moderation, an LLM-based approach was developed to identify rule violations on platforms like Reddit, achieving high true-negative rates but encountering challenges with complex rule interpretation, emphasizing the necessity of human oversight for nuanced cases (Kolla et al., 2024). In behavioral sciences, the LLM-as-a-Judge framework was evaluated for assessing user preferences based on personas, revealing limitations in reliability and consistency due to oversimplified personas, but improved significantly through verbal uncertainty estimation, achieving high agreement with human evaluations for high-certainty cases (Dong et al., 2024). These applications of LLMs as evaluators highlight their growing potential in diverse sectors, emphasizing the need for integrating domain-specific knowledge and refining methodologies.
Moreover, LLMs as evaluators demonstrate significant advantages in qualitative assessments that are difficult to quantify, such as evaluating service quality, analyzing user experience feedback, and assessing creative content like art or literature reviews. LLMs’ capability to understand and generate nuanced language makes them well-suited for subjective evaluation tasks traditionally requiring human judgment. Future research will focus more on these areas, exploring how LLMs as judges can enhance assessment accuracy and consistency where traditional quantitative methods fall short.
@@ 7.1. Reliability
Another significant aspect of reliability is fairness and generalization. Evaluations by LLM-as-a-judge can exhibit considerable inconsistency depending on the context. This is why prompt-based methods are often used to improve LLM-as-a-judge performance. However, challenges related to fairness and generalization may arise due to the sensitivity of prompt engineering. For example, the order of the examples in the context can significantly affect the model’s output, leading to unfair evaluations if the examples are poorly arranged. Moreover, LLMs struggle to handle long context windows effectively, often showing degraded performance or prioritizing later examples in the sequence. These issues raise concerns about fairness and generalization in LLM-based evaluations.
@@ 7.2. Robustness
7.3. Powerful Backbone Model
@@ 8.1. More Reliable LLM-as-a-Judge
Finally, the uncertain and evolving nature of robustness risks underscores the necessity of proactive mitigation strategies. These strategies should include the development of adversarial training techniques tailored to judgment tasks, the integration of robust uncertainty quantification methods, and the implementation of human-in-the-loop systems to oversee critical decisions. By addressing these challenges, we can build more resilient and dependable systems capable of maintaining high levels of reliability even under adversarial conditions.
@@ 8.2. LLM-as-a-Judge for Data Annotation
In contrast, LLM-as-a-judge is a general technique where you use LLM to approximate human labeling. When you ask an LLM to assess qualities like ”faithfulness to source,” ”correctness,” or ”helpfulness,” you define what these terms mean in the evaluation prompt and rely on the semantic relationships the LLM learned from training data. Despite its wide applications, data annotation poses significant challenges for current machine-learning models due to the complexity, subjectivity, and diversity of data. This process requires domain expertise and is resource-intensive, particularly when manually labeling large datasets. Advanced LLMs such as GPT-4 (OpenAI, 2023a), Gemini (Google, 2023), and LLaMA-2 (Touvron et al., 2023c) offer a promising opportunity to revolutionize data annotation. LLMs serve as more than just tools but play a crucial role in improving the effectiveness and precision of data annotation. Their ability to automate annotation tasks (Zhang et al., 2023d), ensure consistency across large volumes of data, and adapt through fine-tuning or prompting for specific domains (Song et al., 2024a; Ma et al., 2024), significantly mitigates the challenges encountered with traditional annotation methods, setting a new standard for what is achievable in the realm of NLP.
Whether in the field of scientific research or industry, we are all still suffering from insufficient target data and domain-specific data, or situations where the data quality is not high enough. Assuming that LLM-as-a-judge can achieve stable performance and be fair and reliable, we can use LLM to annotate data in scenarios where data is insufficient to expand the data. In scenarios with low data quality, we can assess the data quality through LLM, and label the quality tags to achieve the goal of selecting high-quality data. Currently, we have not been able to experimentally rely solely on LLM for a reliable evaluation of various scenarios of data; most of the time, we still rely on human annotation to ensure professionalism and reliability. LLM-as-a-judge often needs to learn from human annotations in order to perform certain labeling tasks.
@@ 8.3. MLLM-as-a-Judge
AI systems are evolving into highly versatile and multifunctional entities. Traditionally, specialized models were required for distinct language processing tasks, such as sentiment analysis, syntactic parsing, and dialogue modeling. However, large language models (LLMs) have demonstrated competence across these tasks using a single set of weights (Srivastava et al., 2022). Similarly, advancements are being made toward unified systems capable of processing multiple data modalities. Instead of employing distinct architectures for processing text, audio, and images, recent models like GPT-4o (OpenAI, 2023b), Gemini (Google, 2023), and LLaVA (Liu et al., 2023b) integrate these capabilities within a single framework. These developments highlight a growing trend toward unification in the structure and functionality of AI systems, which extends to the emerging paradigm of LLM-as-a-Judge.
Currently, MLLM-as-a-Judge frameworks (Chen et al., 2024c) are emerging for evaluating models. However, research exploring how MLLM-as-a-Judge could be applied to the evaluation of data or agents remains limited. Beyond model evaluation, MLLM-as-a-Judge, much like LLM-as-a-Judge, is envisioned to have the capability to assess or annotate data, function as a Reward Model, or serve as a Verifier within intermediate reasoning processes. These expanded roles would allow MLLM-as-a-Judge to contribute more broadly to the AI pipeline.
The future of evaluation lies in developing robust multi-modal evaluators capable of reasoning and assessing complex content spanning text, audio, images, and video. While current multi-modal LLMs exhibit promising capabilities, they often lack the reasoning depth and reliability of their text-based counterparts. Future research must address these limitations, with a focus on enhancing reasoning capabilities, improving reliability, and enabling seamless integration across modalities. A practical multi-modal evaluator has the potential to not only advance AI research but also enable new applications in areas such as multi-modal content moderation and automated knowledge extraction.
@@ 8.5. LLM-as-a-Judge for LLM Optimization
Despite these encouraging developments, current research efforts are still in their infancy. Future work should focus on broadening the application domains and strategies for implementing LLM-as-a-Judge, especially in complex, multi-modal scenarios. Furthermore, a systematic assessment of its reliability and generalization capabilities will be critical for fully realizing the potential of LLM-as-a-Judge in enhancing model performance and robustness.
@@ 9. Conclusion
LLM-as-a-Judge represents an LLM-based paradigm for evaluation, offering a compelling alternative to traditional methods that rely on either human experts or quantitative assessments. By leveraging the powerful capabilities of LLMs, this framework excels in tasks such as scoring, rating, and evaluation, offering exceptional scalability, adaptability, and precision. These qualities align directly with the growing demand for efficient and robust evaluation systems across diverse fields.
However, fully realizing the potential of LLM-as-a-Judge requires addressing significant challenges, particularly those concerning reliability. The evaluation process must account for factors such as consistency, bias mitigation, and contextual adaptability, ensuring that the system operates under well-defined rules while delivering human-aligned and objective results. Each stage of development—from dataset curation and model fine-tuning to evaluation protocol standardization—demands careful consideration to uphold reliability. Simply deploying LLM-as-a-Judge does not guarantee outcomes that are both accurate and aligned with evaluation needs. Building reliable systems requires systematic efforts to evaluate and refine their outputs, while addressing gaps in definitions, practices, and research.
This survey has provided a comprehensive review of the LLM-as-a-Judge landscape, offering insights into its current state and practical strategies for improvement. By addressing these challenges with robust methodologies and innovative benchmarks, LLM-as-a-Judge can evolve into a highly reliable and practical tool for diverse applications. Ultimately, while the current capabilities of LLMs are impressive, their effectiveness as evaluators depends on thoughtful design and systematic refinement. We hope this work serves as a foundation for further research, guiding the development of more reliable LLM-as-a-Judge systems and enabling their broader deployment across academic, industrial, and societal domains.